# Traffic Engineering in Large-scale Networks with Generalizable Graph Neural Networks

### v1

@@ V-D4 Generalization to Unseen Demand Distributions
Remarks: Based on the results, we believe TELGEN can be particularly beneficial in two scenarios: (1) A network operator of a large-scale network can train and evaluate a TELGEN model on smaller subnetworks, such as within a region, and then generalize the pre-trained model to the entire WAN without expensive re-training. This can enable very efficient training than using the entire network topology for training. (2) A networking software provider can train general TE models using small known or random topologies, and ship these models as part of their software to operators. The operators need zero or minimal fine-tuning to generalize these models onto their own, heterogeneous large-scale networks.
VI Conclusions
This paper introduced an innovative methodology for designing customized graph deep learning models and learning algorithms, to tackle the challenges of large-scale network TE problems and ensure generalization to diverse network configurations. TELGEN, a learning-based TE framework, was designed to efficiently solve large-scale TE problems with generalizability. By capturing the relationships among problem variables, objectives, and constraints, TELGEN makes predictions on the TE solution based on a structured graph representation of each problem instance. Two key components for enabled TELGEN’s strong generalization include: flexible model architecture that allows training on diverse topologies, and strong supervision from the classical IPM algorithm for learning instance-agnostic optimization processes.
To evaluate its performance, we trained and tested TELGEN on large-scale randomly generated and real-world network topologies. Extensive experiments demonstrated that TELGEN can solve large-scale TE problems close to optimality, improves prediction time by multiple folds compared to the optimal TE solver, significantly reduces training time on large networks over state-of-the-art learning-based solutions. More importantly, TELGEN generalizes seamlessly to unseen large-scale networks with pre-training on much smaller networks and synthetic traffic demands. With its superior performance and generalizability, TELGEN has the potential to serve as a fundamental building block towards a general-purpose, fully automated software suite for managing large-scale computer networks.

### v2

@@ V-D4 Generalization to Unseen Demand Distributions
Summarizing the above results, we believe TELGEN can be particularly beneficial in two scenarios: (1) A large-scale network operator can train the model on smaller subnetworks, such as within a region, while generalizing the trained model to the entire WAN without expensive re-training. (2) A network software provider may train general TE models using small known or random topologies, and ship these models as part of their software to operators who only need slight fine-tuning to generalize these models onto their own, large-scale networks.
VI Discussion and limitations
In this work, we proposed TELGEN, a novel GNN-based framework for Traffic Engineering that achieves unprecedented generalizability and real-time performance. By learning the algorithmic process of IPM, TELGEN can be trained on small networks and deployed on topologies up to $20\times$ larger with near-optimal results. This addresses the critical scalability and deployment challenges faced by existing learning-based TE methods. While our results are promising, it is important to discuss the limitations of our current work and the future research directions they inspire.
Network Model Applicability: Our work focuses on large-scale wired networks, such as data centers and cloud WANs, and on wireless backbones with point-to-point links, like those in LEO satellite constellations. However, it is not directly applicable to general wireless networks, like cellular or Wi-Fi, where link capacity is dynamic and affected by interference. To address the limitations in network scope, future work could explore new GNN architectures that model the dynamics of general wireless networks.
Dynamic Networks and Online Adaptation: Extending TELGEN to explicitly handle dynamic networks is crucial for practical deployments. Our current model already provides a strong foundation: its generalizability across diverse topologies lets it adapt to discrete topological changes (e.g., link failures) without retraining, and its millisecond-level prediction speed enables real-time reactions to rapidly changing traffic. At present we treat each time snapshot independently. Moving forward, we will incorporate temporal information and online adaptation mechanisms to leverage the continuous nature of network dynamics. These additions aim at a fully adaptive TE solution that proactively captures dynamic network conditions while preserving fast, stable inference.
Distributed and Dynamic Design: TELGEN currently operates in a centralized setting, assuming a global view of the network. It also treats each TE problem as an independent, static snapshot in time. This design does not explicitly model the continuous evolution of traffic demands and may face scalability or single-point-of-failure challenges in massive, decentralized environments. A key future direction is to develop a fully dynamic and distributed TELGEN. This enhanced model would use continuous temporal changes as features to make proactive decisions and employ decentralized agents that coordinate via other novel GNN architectures [49, 48, 18], removing the need for a central controller.
VII Conclusions
This paper proposed a method of utilizing machine learning to tackle the challenges of large-scale network TE problems and generalization to diverse network configurations. TELGEN, a learning-based TE framework is designed to efficiently solve large-scale TE problems with generalizability. Our framework employs modeling TE problems with general graphs, designs a novel double-looped GNN architecture to enable algorithmic alignment, and learns to generalize to unseen network scenarios by mimicking the step-by-step solving process of an optimal TE algorithm. By capturing the relationships among problem variables, objectives, and constraints, TELGEN builds a structured graph. The framework is flexible, allowing for different configurations between training and testing networks by learning the intermediate outcomes of the interior-point method and maintaining input size and topology independence. To evaluate its performance, we trained and tested TELGEN on large-scale randomly generated and real-world network topologies. We demonstrated that TELGEN is suitable and generalizable for large-scale random and practical datasets by using different training and testing dataset configurations. Extensive experiments have demonstrated that TELGEN can solve large-scale TE problems close to optimality with multi-fold improvement in prediction time compared to the optimal TE solver, reduce training time by one orders of magnitude over state-of-the-art learning-based TE solution, and generalize to unseen demand distributions and networks that are much larger than the training networks. Our method outperformed the state-of-the-art learning-based method by prediction gap as well as efficiency. It also outperformed traditional optimization method by efficiency in both small-scale and large-scale datasets. Our proposed TELGEN resolved the scalability issue as well as the generalization challenge for wide-area network traffic engineering.